# Semiparametric Double Reinforcement Learning with Applications to Long-Term Causal Inference

### v3

@@ F.1 Proof of Theorem 8 on uniqueness of Bellman projection
(Existence) By Condition A1, the operator $\mathcal{K}_{P}$ is compact on $\overline{H}_{P}$, so $\mathcal{T}_{P}=I_{P}-\gamma\mathcal{K}_{P}$ is a Fredholm operator of index zero on the Banach space $(\overline{H}_{P},\|\cdot\|_{P})$ (Riesz–Schauder theory, Conway (1994)). In particular, its range $\mathcal{R}:=\mathcal{T}_{P}(\overline{H}_{P})$ is a closed subspace of $L^{2}(P_{A_{0},S_{0}})$.
Now write $Y_{0}=\mu_{P}(S_{0},A_{0})+\epsilon$, where $E_{P}[\epsilon\mid A_{0},S_{0}]=0$. Then for any $q\in\overline{H}_{P}$,
$E_{P}\left[(Y_{0}-\mathcal{T}_{P}(q))^{2}\right]=E_{P}[\epsilon^{2}]+\|\mu_{P}-\mathcal{T}_{P}(q)\|_{P}^{2}.$
Hence the minimizer in (8) corresponds to the projection $\mu_{P,H}:=\operatorname*{argmin}_{\mu\in\mathcal{R}}\|\mu_{P}-\mu\|_{P}^{2}$, which satisfies $\mu_{P,H}\in\mathcal{R}$. Therefore, there exists $q_{P,H}\in\overline{H}_{P}$ such that $\mathcal{T}_{P}(q_{P,H})=\mu_{P,H}$, i.e., a solution to (8) exists. Here, we used the standard fact that the projection onto a closed subspace of a Hilbert space always exists and is unique.
(Uniqueness) Under Condition A2, $\gamma^{-1}$ is not an eigenvalue of $\mathcal{K}_{P}$, so $\mathcal{T}_{P}=I_{P}-\gamma\mathcal{K}_{P}$ has a trivial null space on $\overline{H}_{P}$. Since $\mathcal{T}_{P}$ is Fredholm of index zero (Conway, 1994), injectivity implies surjectivity. Thus, $\mathcal{T}_{P}$ is bijective with a bounded inverse $\mathcal{T}_{P}^{-1}:\mathcal{R}\to\overline{H}_{P}$, and $q_{P,H}=\mathcal{T}_{P}^{-1}(\mu_{P,H})$ is the unique solution to (8). ∎
F.2 Derivation of EIF in Theorem 1 and Theorem 2
@@
By Theorem 8, we have $q_{P,H}=\mathcal{T}_{P}^{-1}(\mu_{P,H})$, where $\mathcal{T}_{P}^{-1}$ is a bounded linear operator and $\mu_{P,H}:=\operatorname*{argmin}_{\mu\in\mathcal{T}_{P}(\overline{H}_{P})}\|\mu_{P}-\mu\|_{P}$ is the projection of $\mu_{P}$ onto the range of $\mathcal{T}_{P}$. Since $\mathcal{T}_{P}$ is invertible on $L^{2}(P)$, its range is closed, and $\overline{H}_{P}$ denotes the closure of $H$ in $L^{2}(P)$. Consequently, $q_{P,H}$ is uniquely identified as an element of the $L^{2}(P)$ closure of $H$.
@@
where $\varphi_{P}$ denotes the efficient influence function (EIF) of $\Psi_{H}$. To compute the pathwise derivative of $\Psi_{H}$, we will use the representation $\Psi_{H}(P)=\langle\mathcal{T}_{P}(\alpha_{P,H}),\mathcal{T}_{P}(q_{P,H})\rangle_{P}=\langle\mathcal{T}_{P}(\alpha_{P,H}),\Pi_{P}(\mu_{P})\rangle_{P}$, which is guaranteed by C1.
@@
$\langle\mathcal{T}_{P}(h),\mu_{P}-\mathcal{T}_{P}(q_{P,H})\rangle_{P}=0\text{ for all }h\in\mathcal{H}.$
@@
$\quad-\frac{d}{d\varepsilon}\langle\mathcal{T}_{P}(h),\mathcal{T}_{P}(q_{P,H})\rangle_{P}\Big{|}_{\varepsilon=0}-\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(h),\mathcal{T}_{P}(q_{P,H})\rangle_{P}\Big{|}_{\varepsilon=0}$
@@
$\quad-\langle\phi,\mathcal{T}_{P}(h)\mathcal{T}_{P}(q_{P,H})-E_{P}[\mathcal{T}_{P}(h)\mathcal{T}_{P}(q_{P,H})]\rangle_{P}-\langle d\mathcal{T}_{P}(h)(\phi),\mathcal{T}_{P}(q_{P,H})\rangle_{P}$
$=\langle d\mathcal{T}_{P}(h)(\phi),\mu_{P}-\mathcal{T}_{P}(q_{P,H})\rangle_{P}+\langle\phi,\mathcal{T}_{P}(h)\{Y_{0}-\mathcal{T}_{P}(q_{P,H})\}\rangle_{P}$
$\quad+\langle\phi,E_{P}[\mathcal{T}_{P}(h)\mathcal{T}_{P}(q_{P,H})]-E_{P}[\mathcal{T}_{P}(h)Y_{0}]\rangle_{P}.$
@@
$\frac{d}{d\varepsilon}\langle\mathcal{T}_{P}(\alpha_{P,H}),\mathcal{T}_{P_{\varepsilon,\phi}}(q_{P_{\varepsilon,\phi}})\rangle_{P}\Big{|}_{\varepsilon=0}$ $=\langle d\mathcal{T}_{P}(\alpha_{P,H})(\phi),\mu_{P}-\mathcal{T}_{P}(q_{P,H})\rangle_{P}+\langle\phi,\mathcal{T}_{P}(\alpha_{P,H})\{Y_{0}-\mathcal{T}_{P}(q_{P,H})\}\rangle_{P}$
$\quad+\langle\phi,E_{P}[\mathcal{T}_{P}(\alpha_{P,H})\mathcal{T}_{P}(q_{P,H})]-E_{P}[\mathcal{T}_{P}(\alpha_{P,H})Y_{0}]\rangle_{P}$
$=\langle d\mathcal{T}_{P}(\alpha_{P,H})(\phi),\mu_{P}-\mathcal{T}_{P}(q_{P,H})\rangle_{P}+\langle\phi,\mathcal{T}_{P}(\alpha_{P,H})\{Y_{0}-\mathcal{T}_{P}(q_{P,H})\}\rangle_{P}$
$=\langle\phi\{\alpha_{P,H}-\gamma\overline{V}_{\alpha_{P,H}}^{\pi}-\mathcal{T}_{P}(\alpha_{P,H})\},\mu_{P}-\mathcal{T}_{P}(q_{P,H})\rangle_{P}$
$\quad+\langle\phi,\mathcal{T}_{P}(\alpha_{P,H})\{Y_{0}-\mu_{P}\}\rangle_{P}.$
@@
$\varphi_{1,P}:w\mapsto\{\alpha_{P,H}(a,s)-\gamma V_{\alpha_{P,H}}^{\pi}(s^{\prime})\}\{\mu_{P}(a,s)-\mathcal{T}_{P}(q_{P,H})(a,s)\}+\mathcal{T}_{P}(\alpha_{P,H})(a,s)\{y-\mu_{P}(a,s)\}.$
@@
$\quad+\frac{d}{d\varepsilon}\langle\mathcal{T}_{P}(\alpha_{P,H}),\mathcal{T}_{P}(q_{P,H})\rangle_{P_{\varepsilon,\phi}}\Big{|}_{\varepsilon=0}.$
@@
$\langle\mathcal{T}_{P}(\alpha_{P,H}),d\mathcal{T}_{P}(q_{P,H})(\phi)\rangle_{P}$ $=-E_{P}\left[\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\left\{\gamma V_{q_{P,H}}^{\pi}(S_{1})+\mathcal{T}_{P}(q_{P,H})-q_{P,H}(A_{0},S_{0})\right\}\phi(Z)\right].$
@@
$\varphi_{P,2}:w\mapsto\mathcal{T}_{P}(\alpha_{P,H})(a,s)\left\{\gamma V_{q_{P,H}}(a,s^{\prime})+\mathcal{T}_{P}(q_{P,H})(a,s)-q_{P,H}(a,s)\right\}\,+\,m(s,a,q_{P,H})-\Psi_{H}(P).$
@@
$w\mapsto$ $\mathcal{T}_{P}(\alpha_{P,H})(a,s)\{y-\mu_{P}(a,s)\}+\{\alpha_{P,H}(a,s)-\gamma V_{\alpha_{P,H}}^{\pi}(a,s^{\prime})\}\{\mu_{P}(a,s)-\mathcal{T}_{P}(q_{P,H})(a,s)\}$
$\quad+\mathcal{T}_{P}(\alpha_{P,H})(a,s)\left\{\gamma V_{q_{P,H}}(a,s^{\prime})+\mathcal{T}_{P}(q_{P,H})(a,s)-q_{P,H}(a,s)\right\}$
$\quad+m(s,a,q_{P,H})-\Psi_{H}(P).$
Assuming a correct model ($\mathcal{T}_{P}(q_{P,H})=\mu_{P}$), it simplifies to:
$\varphi_{P}:w\mapsto$ $\mathcal{T}_{P}(\alpha_{P,H})(a,s)\{y+\gamma V_{q_{P,H}}^{\pi}(a,s^{\prime})-q_{P,H}(a,s)\}$
$\quad+m(s,a,q_{P,H})-\Psi_{H}(P).$
F.3 Derivation of von Mises expansion in Theorem 1 and Theorem 2
We establish the following generalization of the von Mises expansion in Theorem 2, which does not require the nuisance components in the influence function to be compatible with any single distribution $P\in\mathcal{P}$. Let $\widehat{q}_{H},\widehat{\alpha}_{H}\in H$, $\widehat{\mu}_{P}\in L^{\infty}(\lambda)$, and let $\widehat{\mathcal{T}}:L^{\infty}(\lambda)\rightarrow L^{\infty}(\lambda)$ be an arbitrary map. Let $\widehat{P}\in\mathcal{P}$ be a distribution such that $q_{\widehat{P}}^{\pi}=\widehat{q}_{H}$, so that $\Psi(\widehat{P}):=E_{\widehat{P}}[m(S_{0},A_{0},\widehat{q}_{H})]$. Define $\widehat{\varphi}^{*}_{H}$ as the function
@@ F.3 Derivation of von Mises expansion in Theorem 1 and Theorem 2
+$+\left\{\widehat{\alpha}_{H}(a,s)-\gamma V^{\pi}(\widehat{\alpha}_{H})(s^{\prime})-\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(a,s)\right\}\left\{\mu_{0}(a,s)-\widehat{\mathcal{T}}(\widehat{q}_{H})(a,s)\right\}$
$+m(s,a,\widehat{q}_{H})-\Psi(\widehat{P}).$
@@
$\quad+P_{0}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})-\mathcal{T}_{0}(\widehat{\alpha}_{H})\right\}(\mu_{0}-\widehat{\mu})\right]$
$\quad+P_{0}\left[\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})-\mathcal{T}_{0}(\widehat{\alpha}_{H})\}(\widehat{\mathcal{T}}(\widehat{q}_{H})-\mathcal{T}_{0}(\widehat{q}_{H}))\right].$
@@
Let $\widehat{P}$ be a distribution compatible with $\widehat{q}_{H}$ and the marginal distribution of $(S_{0},A_{0})$ used to compute the term $\Psi_{H}(\widehat{P})$ in $\widehat{\varphi}^{*}$. By the law of iterated expectations, it holds that
@@
$=E_{P}\left[\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})\{\mu_{P}(A_{0},S_{0})-\widehat{\mu}(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\{\widehat{\alpha}_{H}(A_{0},S_{0})-\gamma V^{\pi}(\widehat{\alpha}_{H})(S_{1})\}\{\widehat{\mu}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})\left\{\gamma V^{\pi}(\widehat{q}_{H})(S_{1})+\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\widehat{q}_{H}(A_{0},S_{0})\right\}\right]$
$\quad+E_{P}\left[m(S_{0},A_{0},\widehat{q}_{H})-m(S_{0},A_{0},q_{P,H})\right].$
@@
$=E_{P}\left[\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\left\{\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})-\mu_{P}(A_{0},S_{0})\right\}\right],$
where we used that $\mathcal{T}_{P}(q_{P,H})$ is the $L^{2}(P)$ projection of $\mu_{P}$ onto $\mathcal{T}_{P}(\overline{H}_{P})$. In addition, applying the law of iterated expectations applied to the second and third terms, we find:
$\widehat{R}_{H}^{*}(P)$ $=E_{P}\left[\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})\{\mu_{P}(A_{0},S_{0})-\widehat{\mu}(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\{\widehat{\mu}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})\right\}\right]$
$\quad+E_{P}\left[\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\left\{\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})-\mu_{P}(A_{0},S_{0})\right\}\right].$
@@
$\widehat{R}_{H}^{*}(P)$ $=E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\right\}\{\mu_{P}(A_{0},S_{0})-\widehat{\mu}(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\{\mu_{P}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})\right\}\right]$
$\quad+E_{P}\left[\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\left\{\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})-\mu_{P}(A_{0},S_{0})\right\}\right].$
@@
$\widehat{R}_{H}^{*}(P)$ $=E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\right\}\{\mu_{P}(A_{0},S_{0})-\widehat{\mu}(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\{\mu_{P}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})\right\}\right]$
$\quad+E_{P}\left[\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\mu_{P}(A_{0},S_{0})\right\}\right].$

### v4

@@ H.1 Proof of Theorem 7 on uniqueness of Bellman projection
(Existence) By Condition A1, the operator $\mathcal{K}_{P}$ is compact on $\overline{H}_{P}$, so $\mathcal{T}_{P}=I_{P}-\gamma\mathcal{K}_{P}$ is a Fredholm operator of index zero on the Banach space $(\overline{H}_{P},\|\cdot\|_{P})$ (Riesz–Schauder theory, conway1994course). In particular, its range $\mathcal{R}:=\mathcal{T}_{P}(\overline{H}_{P})$ is a closed subspace of $L^{2}(P_{A_{0},S_{0}})$.
Now write $Y_{0}=r_{P}(S_{0},A_{0})+\epsilon$, where $E_{P}[\epsilon\mid A_{0},S_{0}]=0$. Then for any $q\in\overline{H}_{P}$,
$E_{P}\left[(Y_{0}-\mathcal{T}_{P}(q))^{2}\right]=E_{P}[\epsilon^{2}]+\|r_{P}-\mathcal{T}_{P}(q)\|_{P}^{2}.$
Hence the minimizer in (7) corresponds to the projection $r_{P,H}:=\operatorname*{argmin}_{r\in\mathcal{R}}\|r_{P}-r\|_{P}^{2}$, which satisfies $r_{P,H}\in\mathcal{R}$. Therefore, there exists $q_{P,H}\in\overline{H}_{P}$ such that $\mathcal{T}_{P}(q_{P,H})=r_{P,H}$, i.e., a solution to (7) exists. Here, we used the standard fact that the projection onto a closed subspace of a Hilbert space always exists and is unique.
(Uniqueness) Under Condition A2, $\gamma^{-1}$ is not an eigenvalue of $\mathcal{K}_{P}$, so $\mathcal{T}_{P}=I_{P}-\gamma\mathcal{K}_{P}$ has a trivial null space on $\overline{H}_{P}$. Since $\mathcal{T}_{P}$ is Fredholm of index zero (conway1994course), injectivity implies surjectivity. Thus, $\mathcal{T}_{P}$ is bijective with a bounded inverse $\mathcal{T}_{P}^{-1}:\mathcal{R}\to\overline{H}_{P}$, and $q_{P,H}=\mathcal{T}_{P}^{-1}(r_{P,H})$ is the unique solution to (7). ∎
H.2 Derivation of EIF in Theorem 1
@@
By Theorem 7, we have $q_{P,H}=\mathcal{T}_{P}^{-1}(r_{P,H})$, where $\mathcal{T}_{P}^{-1}$ is a bounded linear operator and $r_{P,H}:=\operatorname*{argmin}_{r\in\mathcal{T}_{P}(\overline{H}_{P})}\|r_{P}-r\|_{P}$ is the projection of $r_{P}$ onto the range of $\mathcal{T}_{P}$. Since $\mathcal{T}_{P}$ is invertible on $L^{2}(P)$, its range is closed, and $\overline{H}_{P}$ denotes the closure of $H$ in $L^{2}(P)$. Consequently, $q_{P,H}$ is uniquely identified as an element of the $L^{2}(P)$ closure of $H$.
@@
where $\varphi_{P}$ denotes the efficient influence function (EIF) of $\Psi_{H}$. To compute the pathwise derivative of $\Psi_{H}$, we will use the representation $\Psi_{H}(P)=\langle\mathcal{T}_{P}(\alpha_{P,H}),\mathcal{T}_{P}(q_{P,H})\rangle_{P}=\langle\mathcal{T}_{P}(\alpha_{P,H}),\Pi_{P}(r_{P})\rangle_{P}$, which is guaranteed by C1.
@@
$\langle\mathcal{T}_{P}(h),r_{P}-\mathcal{T}_{P}(q_{P,H})\rangle_{P}=0\text{ for all }h\in\mathcal{H}.$
@@
$\qquad-\frac{d}{d\varepsilon}\langle\mathcal{T}_{P}(h),\mathcal{T}_{P}(q_{P,H})\rangle_{P}\Big|_{\varepsilon=0}-\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(h),\mathcal{T}_{P}(q_{P,H})\rangle_{P}\Big|_{\varepsilon=0}$
@@
$\qquad-\langle\phi,\mathcal{T}_{P}(h)\mathcal{T}_{P}(q_{P,H})-E_{P}[\mathcal{T}_{P}(h)\mathcal{T}_{P}(q_{P,H})]\rangle_{P}-\langle d\mathcal{T}_{P}(h)(\phi),\mathcal{T}_{P}(q_{P,H})\rangle_{P}$
$=\langle d\mathcal{T}_{P}(h)(\phi),r_{P}-\mathcal{T}_{P}(q_{P,H})\rangle_{P}+\langle\phi,\mathcal{T}_{P}(h)\{Y_{0}-\mathcal{T}_{P}(q_{P,H})\}\rangle_{P}$
$\qquad+\langle\phi,E_{P}[\mathcal{T}_{P}(h)\mathcal{T}_{P}(q_{P,H})]-E_{P}[\mathcal{T}_{P}(h)Y_{0}]\rangle_{P}.$
@@
$\frac{d}{d\varepsilon}\langle\mathcal{T}_{P}(\alpha_{P,H}),\mathcal{T}_{P_{\varepsilon,\phi}}(q_{P_{\varepsilon,\phi}})\rangle_{P}\Big|_{\varepsilon=0}$ $=\langle d\mathcal{T}_{P}(\alpha_{P,H})(\phi),r_{P}-\mathcal{T}_{P}(q_{P,H})\rangle_{P}+\langle\phi,\mathcal{T}_{P}(\alpha_{P,H})\{Y_{0}-\mathcal{T}_{P}(q_{P,H})\}\rangle_{P}$
$\qquad+\langle\phi,E_{P}[\mathcal{T}_{P}(\alpha_{P,H})\mathcal{T}_{P}(q_{P,H})]-E_{P}[\mathcal{T}_{P}(\alpha_{P,H})Y_{0}]\rangle_{P}$
$=\langle d\mathcal{T}_{P}(\alpha_{P,H})(\phi),r_{P}-\mathcal{T}_{P}(q_{P,H})\rangle_{P}+\langle\phi,\mathcal{T}_{P}(\alpha_{P,H})\{Y_{0}-\mathcal{T}_{P}(q_{P,H})\}\rangle_{P}$
$=\langle\phi\{\alpha_{P,H}-\gamma\overline{V}_{\alpha_{P,H}}^{\pi}-\mathcal{T}_{P}(\alpha_{P,H})\},r_{P}-\mathcal{T}_{P}(q_{P,H})\rangle_{P}$
$\qquad+\langle\phi,\mathcal{T}_{P}(\alpha_{P,H})\{Y_{0}-r_{P}\}\rangle_{P}.$
@@
$\varphi_{1,P}:w\mapsto\{\alpha_{P,H}(a,s)-\gamma V^{\pi}(\alpha_{P,H})(s^{\prime})\}\{r_{P}(a,s)-\mathcal{T}_{P}(q_{P,H})(a,s)\}+\mathcal{T}_{P}(\alpha_{P,H})(a,s)\{y-r_{P}(a,s)\}.$
@@
$\qquad+\frac{d}{d\varepsilon}\langle\mathcal{T}_{P}(\alpha_{P,H}),\mathcal{T}_{P}(q_{P,H})\rangle_{P_{\varepsilon,\phi}}\Big|_{\varepsilon=0}.$
@@
$\langle\mathcal{T}_{P}(\alpha_{P,H}),d\mathcal{T}_{P}(q_{P,H})(\phi)\rangle_{P}$ $=-E_{P}\left[\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\left\{\gamma V^{\pi}(q_{P,H})(S_{1})+\mathcal{T}_{P}(q_{P,H})-q_{P,H}(A_{0},S_{0})\right\}\phi(Z)\right].$
@@
$\varphi_{P,2}:w\mapsto\mathcal{T}_{P}(\alpha_{P,H})(a,s)\left\{\gamma V^{\pi}(q_{P,H})(a,s^{\prime})+\mathcal{T}_{P}(q_{P,H})(a,s)-q_{P,H}(a,s)\right\}\,+\,m(s,a,q_{P,H})-\Psi_{H}(P).$
@@
$w\mapsto$ $\mathcal{T}_{P}(\alpha_{P,H})(a,s)\{y-r_{P}(a,s)\}+\{\alpha_{P,H}(a,s)-\gamma V^{\pi}(\alpha_{P,H})(a,s^{\prime})\}\{r_{P}(a,s)-\mathcal{T}_{P}(q_{P,H})(a,s)\}$
$\qquad+\mathcal{T}_{P}(\alpha_{P,H})(a,s)\left\{\gamma V^{\pi}(q_{P,H})(a,s^{\prime})+\mathcal{T}_{P}(q_{P,H})(a,s)-q_{P,H}(a,s)\right\}$
$\qquad+m(s,a,q_{P,H})-\Psi_{H}(P).$
Assuming a correct model ($\mathcal{T}_{P}(q_{P,H})=r_{P}$), it simplifies to:
$\varphi_{P}:w\mapsto$ $\mathcal{T}_{P}(\alpha_{P,H})(a,s)\{y+\gamma V^{\pi}(q_{P,H})(a,s^{\prime})-q_{P,H}(a,s)\}$
$\qquad+m(s,a,q_{P,H})-\Psi_{H}(P).$
H.3 Derivation of von Mises expansion in Theorem 1
We establish the following generalization of the von Mises expansion in Theorem 1, which does not require the nuisance components in the influence function to be compatible with any single distribution $P\in\mathcal{P}$. Let $\widehat{q}_{H},\widehat{\alpha}_{H}\in H$, $\widehat{r}_{P}\in L^{\infty}(\lambda)$, and let $\widehat{\mathcal{T}}:L^{\infty}(\lambda)\rightarrow L^{\infty}(\lambda)$ be an arbitrary map. Let $\widehat{P}\in\mathcal{P}$ be a distribution such that $q_{\widehat{P}}^{\pi}=\widehat{q}_{H}$, so that $\Psi(\widehat{P}):=E_{\widehat{P}}[m(S_{0},A_{0},\widehat{q}_{H})]$. Define $\widehat{\varphi}^{*}_{H}$ as the function
@@ H.3 Derivation of von Mises expansion in Theorem 1
$+\left\{\widehat{\alpha}_{H}(a,s)-\gamma V^{\pi}(\widehat{\alpha}_{H})(s^{\prime})-\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(a,s)\right\}\left\{r_{0}(a,s)-\widehat{\mathcal{T}}(\widehat{q}_{H})(a,s)\right\}$
[…]
@@
[…]
$\qquad+P_{0}\left[\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})-\mathcal{T}_{0}(\widehat{\alpha}_{H})\}(\widehat{\mathcal{T}}(\widehat{q}_{H})-\mathcal{T}_{0}(\widehat{q}_{H}))\right].$
@@
[…]
@@
$=E_{P}\left[\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})\{r_{P}(A_{0},S_{0})-\widehat{r}(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\{\widehat{\alpha}_{H}(A_{0},S_{0})-\gamma V^{\pi}(\widehat{\alpha}_{H})(S_{1})\}\{\widehat{r}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})\left\{\gamma V^{\pi}(\widehat{q}_{H})(S_{1})+\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\widehat{q}_{H}(A_{0},S_{0})\right\}\right]$
$\qquad+E_{P}\left[m(S_{0},A_{0},\widehat{q}_{H})-m(S_{0},A_{0},q_{P,H})\right].$
@@
$=E_{P}\left[\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\left\{\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})-r_{P}(A_{0},S_{0})\right\}\right],$
where we used that $\mathcal{T}_{P}(q_{P,H})$ is the $L^{2}(P)$ projection of $r_{P}$ onto $\mathcal{T}_{P}(\overline{H}_{P})$. In addition, applying the law of iterated expectations applied to the second and third terms, we find:
$\widehat{R}_{H}^{*}(P)$ $=E_{P}\left[\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})\{r_{P}(A_{0},S_{0})-\widehat{r}(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\{\widehat{r}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})\right\}\right]$
$\qquad+E_{P}\left[\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\left\{\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})-r_{P}(A_{0},S_{0})\right\}\right].$
@@
$\widehat{R}_{H}^{*}(P)$ $=E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\right\}\{r_{P}(A_{0},S_{0})-\widehat{r}(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\{r_{P}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})\right\}\right]$
$\qquad+E_{P}\left[\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\left\{\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})-r_{P}(A_{0},S_{0})\right\}\right].$
@@
$\widehat{R}_{H}^{*}(P)$ $=E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\right\}\{r_{P}(A_{0},S_{0})-\widehat{r}(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\{r_{P}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})\right\}\right]$
$\qquad+E_{P}\left[\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-r_{P}(A_{0},S_{0})\right\}\right].$